# GTR: Guided Thought Reinforcement Prevents Thought Collapse in RL-based VLM Agent Training

### v1

@@ 5.2 Process Guidance Improves VLM Decision-making Capabilities
Figure 5: Training curves on the 24 points game environment. Compared to the baseline methods, our GTR framework integrating process guidance with RL achieves better performance while maintaining a rational reasoning process. Curves are smoothed for better readability. Since GTR and SFT-only employ truncation strategies, we plot $\gamma=0.9$ discounted returns in the figure for a fair comparison.
@@
As illustrated in Figure 6 and Table 6, our GTR algorithm demonstrates significant and consistent performance improvements on the Points24 task compared to existing models, e.g., a 3x-5x increase in success rate and a higher return. This achievement underscores the importance of process guidance in RL training for VLMs in challenging tasks. Furthermore, GTR outperforms the thought cloning method and surpasses GPT4o with tool function calling - the corrector model itself. This demonstrates that RL allows the agent to go beyond simple imitation, exploring and discovering superior strategies under the guidance of the corrector.
@@
Table 1: Performance of different models in other tasks in $\mathtt{gym\_cards}$. On simpler tasks where thought collapse is not evident, GTR still achieves improvements over RL4VLM and is comparable to the pretrained model 10x larger in size. SR - success rate, ER - episode return, * - reported in previous work.
We also assess the performance of GTR on the other tasks in $\mathtt{gym\_cards}$, shown in Table 1. Due to smaller state and action spaces and significantly shorter episode lengths, thought collapse is not prominent in these tasks. Nevertheless, GTR still achieves noticeable improvements, performing comparably to Qwen-2-VL, which is 10 times larger than our model in size, highlighting the effectiveness of our framework.
@@
Figure 7: Comparison of training curves between GTR and RL4VLM in the ALFWorld environment. RL4VLM fails to effectively facilitate model learning, leading to thought collapse. GTR, however, enables the agent’s performance to improve steadily, ultimately achieving superior results.
@@
In Figure 8, we compare the training curves of GTR and RL4VLM on the ALFWorld embodied environment. The results show that RL4VLM, which relies solely on outcome rewards, struggles to provide effective guidance, with both win rates and returns declining, exhibiting signs of thought collapse. In contrast, with thought guidance in GTR, although the corrector model may not be as precise as an expert policy, it is sufficient to stimulate the agent’s reasoning ability. Table 8 lists the success rates of different methods on tasks in ALFWorld. It is worth noting that using the textual descriptions of scenes in ALFWorld as inputs significantly reduces the difficulty of recognition and decision-making, proved by the heavy reliance observed in the output thoughts. This creates an unreasonable advantage over our setting. Nevertheless, GTR can still achieve competitive success rates through reinforcement learning.
@@ 5.3 Ablation Study
The necessity of process guidance thoughtout training
We aim to investigate whether adjusting the thought cloning SFT loss and PPO loss weights can enable a training regime where the agent imitates reasoning early on and focuses on free exploration later. To this end, we apply cosine annealing to the weight of the thought cloning loss, gradually reducing its proportion in the total loss function. As shown in Figure 9, the training curve with annealing fails to achieve the same breakthrough as GTR. A closer examination of the model’s outputs reveals that relaxing process guidance may lead to a return to thought collapse, compromising the consistency of reasoning. This fact proves that process guidance is essential throughout the training process to maintain the stability and effectiveness of RL training.
The importance of capable corrector model
We experiment with removing the tool function calling ability from the GPT-4o corrector, which significantly impairs its ability to analyze and solve the 24-point game task. In this setup, the reference thought provided to the agent lacks rationality, and performance did not improve, which aligns with intuition. Although the agent retained some visual capabilities, its reasoning remained illogical and disconnected from action decisions. This demonstrates that the corrector model must possess enough analytical and problem-solving abilities to synergize with RL training effectively. It also highlights the value of incorporating task-specific knowledge when using general-purpose foundation models as correctors.
@@
In Figure 9, we compare the training curves of GTR with and without DAgger. Results show that removing DAgger still allows the model to achieve increased performance in the early stage, but further breakthroughs become increasingly difficult as training progresses. This indicates that distribution shift and forgetting caused by the evolving thought cloning dataset indeed hinder the continual improvement of RL learning. This issue is alleviated by adopting the DAgger method from imitation learning, which constantly expands the dataset for thought cloning.
@@
We also compare thought cloning with performing SFT on both thoughts and actions. Figure 10 shows that SFT on the full response does not yield promising results. From detailed outputs during training, we identified instances where mismatches between thoughts and actions were incorrectly validated due to corrector hallucinations. We conclude that the constraints on actions from the environment and corrector have interference, and the stronger SFT constraints make the agent more susceptible to corrector hallucinations, undermining its ability to adjust based on environmental feedback. Cloning thoughts alone remains a more balanced and robust approach.
Figure 10: Comparing thought cloning with supervise finetuning on both thoughts and actions. This full response cloning approach does not yield satisfactory results due to its vulnerability to the corrector’s hallucinations.
@@ 6 Conclusion
During the RL finetuning of VLM agents for challenging tasks, we identified the thought collapse issue, where the lack of thought supervision leads the agent to generate state-irrelevant reasoning, ultimately losing its ability to think coherently. Therefore, we propose the Guided Thought Reinforcement framework, which introduces an automated corrector to refine the agent’s reasoning on the fly. This simple yet scalable approach provides effective process guidance. Combining thought reinforcement with the well-established RL framework, GTR unleashes the agent’s decision-making capabilities, enabling significantly smaller models to achieve substantial advantages in complex, long-horizon tasks.
While our study suggests that process guidance is effective for RL training of VLM agents in multi-step tasks, we have not extensively explored the approach promoting o1-like multi-round CoT for action sequence reasoning, which remains an interesting direction for future research. Additionally, due to resource limitations, our research primarily focuses on 7b-scale models. Larger models may further enhance agent performance.
@@ Appendix B Additional Details on Training
Drawing inspiration from the RLHF training framework [28] and prior related work [61], we perform 1 epoch of supervised fine-tuning on the base LLaVa-v1.6-mistral-7b model [22, 21, 23] before RL training, which is referred to as LLaVa-sft in the results. The datasets are sourced from the RL4VLM paper [61], with labels for the $\mathtt{gym\_cards}$ environment provided by a task solver and labels for the ALFWorld environment generated by GPT-4V.
@@
Appendix C Prompts of the Corrector Model
@@
Prompt adopted by VLM corrector model for Points24 task System Prompt: You are an expert 24-point card game player. You are observing four cards in the image and the current formula. The goal is to output a formula that evaluates to 24, and each number can only be used once. The number or operator include [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’, ’9’, ’10’, ’+’, ’-’, ’*’, ’/’, ’(’, ’)’, ’=’], and the chosen number or operator will be appended to the current formula to reach the correct target formula that evaluates to 24. Note that ’J’, ’Q’, and ’K’ count as ’10’. Query: You will be given the current formula, the thought of a player playing this game, and a target formula. The player’s thought may be wrong, please evaluate its correctness by the following aspects: (1) What are the four cards in the image? If the target formula is ’NOT DETERMINED’, use the ’find_all_correct_formulas’ tool function to find all possible correct formulas by the four cards in the image. Remember the correct formulas, and do not output the result. (2) What are the recognized card ranks in the thought? According to the rules, does the ranks in the thought match your observation in question (1), regardless of the order? (3) What is the proposed formula the player is trying to reach in the thought? Does the proposed formula match the target formula or, if the target formula is ’NOT DETERMINED’, one of the possible correct formulas in question (1)? (4) Does the player choose the correct action to reach the proposed formula or choose ’=’ if the current formula is complete? Please briefly answer the above questions, then give your final evaluation. If the thought is incorrect, use all available information for thought correction: determine the next single number or character to append to the current formula and finally provide the correct thought. Your response should be a valid json file in the following format: { "answer1": {Text, answer to the first question}, "answer2": {Text, answer to the second question}, "answer3": {Text, answer to the third question}, "answer4": {Text, answer to the third question}, "evaluation": {YES or NO}, "possible_solution": {YES or NO, indicating whether there is a possible solution. None if the thought is correct}, "target_formula": {The given target formula if it is not None. The proposed formula in the thought if the thought is correct. Otherwise, choose an appropriate target formula from all possible correct formulas obtained from the tool function for the player to reach. }, "correction": {Json object, the correct thought. None if the thought is correct} } [Current Formula] ... [Thought] ... [Target Formula] ...
@@
Prompt adopted by VLM corrector model for ALFWorld task System Prompt: You are an expert in the ALFRED Embodied Environment. The environment requires the player to navigate, take certain objects, interact with objects if necessary, and finally put objects in the designated place to complete the task. Query: You will be given the visual observation and thought of a player in this environment. The task is to ... You are also given the previous actions the player has taken: ... All admissible actions of the current situation are: ... Please evaluate if the reasoning is correct by the following aspects: (1) What objects are in your sight and whether you are holding a certain object? Does the thought correctly identify the image? (2) Based on the task description and the action history, what should be the player’s next sub-goal (notice that the tasks require the player to first pick up certain objects, interact with receptacles if the task is cooling, heating, cleaning or looking in light, and finally placing the object)? Does the thought align with the sub-goal? (3) Based on the task description and the action history, does the player choose one of the admissible actions to reach the sub-goal? Does the action take effect? If the target object is not in sight, go to an unexplored location; if there is a required object, take it; if the task requires cooling, heating, cleaning, or looking in light, navigate and interact with the receptacles. Please briefly answer the above questions, then give your final evaluation. If the thought is incorrect, include all available information for thought correction: choose one correct step from the admissible actions for the player to finish the task, and finally provide the correct thought. Your response should be a valid json file in the following format: { "answer1": {Text, answer to the first question}, "answer2": {Text, answer to the second question}, "answer3": {Text, answer to the third question}, "evaluation": {YES or NO}, "correction": {Json object, the correct thought. None if the thought is correct} } [The player’s thought] ...
Appendix D Example Trajectories of Thought Collapse
@@
In this section, we present real episode examples from the Points24 and ALFWorld tasks that exhibit thought collapse during RL training. These examples demonstrate how the model’s outputs degrade, becoming state-irrelevant and irrational, significantly impairing the agent’s decision-making capabilities. This catastrophic phenomenon is the core issue addressed in our work.
Figure 14: Examples of thought collapse trajectories in Points24 and ALFWorld.

### v2

@@ 5.2 Process Guidance Improves VLM Decision-making Capabilities
Figure 4: Training curves on the 24 points game environment. Compared to the baseline methods, our GTR framework, integrating process guidance with RL, achieves better performance while maintaining a rational reasoning process. Curves are smoothed for better readability. Since GTR and SFT-only employ truncation strategies, we plot $\gamma=0.9$ discounted returns in the figure for a fair comparison.
@@
As illustrated in Figure 5 and Table 5, our GTR algorithm demonstrates significant and consistent performance improvements on the Points24 task compared to existing models by a 3- to 5-fold increase in success rate and a higher return. This achievement highlights the significance of process guidance in RL training for VLMs in complex tasks. Furthermore, GTR outperforms the thought cloning method and surpasses GPT4o with tool function calling, the corrector model itself. This demonstrates that RL allows the agent to go beyond simple imitation, exploring and discovering superior strategies under the guidance of the corrector.
Figure 6: Comparison of training curves on Qwen2.5-VL-7B agents. GTR boosts model’s reasoning capability, even reaching o3-level performance.
We also conducted experiments using the more recent and powerful Qwen2.5-VL-7B model. As shown in Figure 6, GTR continues to deliver substantial performance improvements, even enabling the agent to reach o3-level performance within a short training period. In contrast, the RL4VLM’s performance degrades as the training budget increases.
We also assess the performance of GTR on the other tasks in $\mathtt{gym\_cards}$, shown in Table 1. Due to smaller state and action spaces and significantly shorter episode lengths, thought collapse is not prominent in these tasks. Nevertheless, GTR still achieves noticeable improvements, performing comparably to Qwen2-VL, which is 10 times larger than our model in size, highlighting the effectiveness of our framework.
In Figure 8, we compare the training curves of GTR and RL4VLM on the ALFWorld embodied environment. The results show that RL4VLM, which relies solely on outcome rewards, struggles to provide effective guidance, with both win rates and returns declining, exhibiting signs of thought collapse. In contrast, with thought guidance in GTR, although the corrector model may not be as precise as an expert policy, it is sufficient to stimulate the agent’s reasoning ability. Table 8 lists the success rates of different methods on tasks in ALFWorld. It is worth noting that using the textual descriptions of scenes in ALFWorld as inputs significantly reduces the difficulty of recognition and decision-making, as proved by the heavy reliance observed in the output thoughts. This creates an unreasonable advantage over our setting. Nevertheless, GTR can still achieve competitive success rates through reinforcement learning.
@@
Table 1: Performance of different models in other tasks in $\mathtt{gym\_cards}$. On simpler tasks where thought collapse is not evident, GTR still achieves improvements over RL4VLM and is comparable to the pretrained model 10x larger. SR - success rate, ER - episode return, * - reported in previous work.
@@
Figure 7: Comparison of training curves between GTR and RL4VLM in the ALFWorld environment. RL4VLM fails to facilitate model learning, leading to thought collapse effectively. GTR, however, enables the agent’s performance to improve steadily, ultimately achieving superior results.
@@
[…]
Figure 10: Comparing thought cloning with SFT on both thoughts and actions. This full response cloning approach does not yield satisfactory results due to its vulnerability to the corrector’s hallucinations.
@@ 5.3 Ablation Study
The necessity of process guidance throughout training
We aim to investigate whether adjusting the loss weights can enable a training regime where the agent imitates reasoning early on and focuses on free exploration later. To this end, we apply cosine annealing to the weight of the thought cloning loss, gradually reducing its proportion in the total loss function. As shown in Figure 10, the training curve with annealing fails to achieve the same breakthrough as GTR. A closer examination of the model’s outputs reveals that relaxing process guidance may lead to a return to thought collapse. This fact proves that process guidance is essential throughout the training process to maintain the stability and effectiveness of RL training.
The importance of a capable corrector model
We experiment with removing the tool function calling ability from the GPT-4o corrector, which significantly impairs its ability to analyze and solve the 24-point game task. In this setup, the reference thought provided to the agent lacks rationality, and performance did not improve, which aligns with intuition. Although the agent retained some visual capabilities, its reasoning remained illogical and disconnected from action decisions. This demonstrates that the corrector model must possess sufficient analytical and problem-solving abilities to synergize effectively with RL training. It also highlights the value of incorporating task-specific knowledge when using general-purpose foundation models as correctors.
@@
In Figure 10, we compare the training curves of GTR with and without DAgger. Results show that removing DAgger still allows the model to achieve increased performance in the early stage, but further breakthroughs become difficult as training progresses. This indicates that distribution shift and forgetting caused by the evolving thought cloning dataset indeed hinder the continual improvement of RL learning. This issue is alleviated by adopting the DAgger method from imitation learning, which constantly expands the dataset for thought cloning.
@@
We also try performing SFT on both thoughts and actions. Figure 10 shows that SFT on the full response does not yield promising results. From detailed outputs, we find instances where mismatches between thoughts and actions are validated due to corrector hallucinations. We conclude that the constraints on actions from the environment and corrector interfere, and the stronger SFT constraints make the agent more susceptible to corrector hallucinations, undermining its ability to adjust based on environmental feedback. Cloning thoughts alone remains a more balanced and robust approach.
@@ 6 Conclusion
During the RL finetuning of VLM agents for challenging tasks, we identified the thought collapse issue, where the lack of thought supervision leads the agent to generate state-irrelevant reasoning, ultimately losing its ability to think coherently. Therefore, we propose the Guided Thought Reinforcement framework, which introduces an automated corrector to refine the agent’s reasoning on the fly. This simple and scalable approach provides effective process guidance. Combining thought reinforcement with the well-established RL framework, GTR unleashes the agent’s decision-making capabilities, enabling significantly smaller models to achieve substantial advantages in complex, long-horizon tasks.
While our study suggests that process guidance is effective for RL training of VLM agents in multi-step tasks, we have not explored the approach promoting o1-like long CoT for action sequence reasoning, which remains an interesting direction for future research. Additionally, due to resource limitations, our research primarily focuses on 7B-scale models. Larger models may further enhance agent performance.
@@ Appendix B Additional Details on Training
B.1 Training Setting
Drawing inspiration from the RLHF training framework [28] and prior related work [61], we perform one epoch of supervised fine-tuning on the base LLaVA-v1.6-mistral-7B model [22, 21, 23] before RL training, which is referred to as LLaVA-sft in the results. The datasets are sourced from the RL4VLM paper [61], with labels for the $\mathtt{gym\_cards}$ environment provided by a task solver and labels for the ALFWorld environment generated by GPT-4V.
B.2 Hyperparameters
@@
B.3 Computational Overhead
Running a large corrector model at each RL step introduces additional computational overhead during training. To provide a more comprehensive comparison of corrector models of different types and sizes, we evaluate the performance, cost, and training time of both open-source and closed-source models, as well as large and small models. The results in Table 3 show that although open-source models can reduce overhead, the performance of Qwen2.5-VL-72B is undermined by its sub-optimal tool-use capabilities, and the 7B version fails to follow proper correction formats.
Corrector Model Performance Token Usage (Cost) Time
GPT-4o 17.5% 33.5M (~\$463.5) 86h
Qwen2.5-VL-72B 6.5% 33.8M (~\$91.6) 110h
Qwen2.5-VL-7B N/A 31.2M (~\$19.9) 56h
Table 3: Ablation study on computational overhead across different corrector models.
Appendix C Additional results of ALFWorld with textual observation
We observe that RL4VLM’s performance on ALFWorld in its original paper is attributed to precise textual observations, which notably compensates for the agent’s limited visual recognition and reasoning capabilities. This is also the primary reason why we remove the text description in our experiments. Nevertheless, we include the performance of RL4VLM and GTR with the presence of textual observations. The results in Table 4 demonstrate that GTR does not need textual observations but achieves competitive performance as the corrector effectively bridges the gap between vision and text.
Success Rate
RL4VLM w/ text 21.7%
RL4VLM w/o text 5.4%
GTR w/ text 21.0%
GTR w/o text 17.8%
Table 4: Performance comparison of ALFWorld with and without textual observations.
Appendix D Continual Training of GTR
To evaluate the performance of the GTR algorithm over extended training durations, we present results of GTR trained for 30k steps on the Points24 task. As shown in Figure 14, GTR is able to maintain excellent performance.
Figure 14: Training the VLM agent with GTR for 30,000 steps.
Appendix E Prompts of the Corrector Model
@@
Prompt adopted by the VLM corrector model for the Points24 task System Prompt: You are an expert 24-point card game player. You are observing four cards in the image and the current formula. The goal is to output a formula that evaluates to 24, and each number can only be used once. The number or operator include [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’, ’9’, ’10’, ’+’, ’-’, ’*’, ’/’, ’(’, ’)’, ’=’], and the chosen number or operator will be appended to the current formula to reach the correct target formula that evaluates to 24. Note that ’J’, ’Q’, and ’K’ count as ’10’. Query: You will be given the current formula, the thought of a player playing this game, and a target formula. The player’s thought may be wrong, please evaluate its correctness by the following aspects: (1) What are the four cards in the image? If the target formula is ’NOT DETERMINED’, use the ’find_all_correct_formulas’ tool function to find all possible correct formulas by the four cards in the image. Remember the correct formulas, and do not output the result. (2) What are the recognized card ranks in the thought? According to the rules, does the ranks in the thought match your observation in question (1), regardless of the order? (3) What is the proposed formula the player is trying to reach in the thought? Does the proposed formula match the target formula or, if the target formula is ’NOT DETERMINED’, one of the possible correct formulas in question (1)? (4) Does the player choose the correct action to reach the proposed formula or choose ’=’ if the current formula is complete? Please briefly answer the above questions, then give your final evaluation. If the thought is incorrect, use all available information for thought correction: determine the next single number or character to append to the current formula and finally provide the correct thought. Your response should be a valid json file in the following format: { "answer1": {Text, answer to the first question}, "answer2": {Text, answer to the second question}, "answer3": {Text, answer to the third question}, "answer4": {Text, answer to the third question}, "evaluation": {YES or NO}, "possible_solution": {YES or NO, indicating whether there is a possible solution. None if the thought is correct}, "target_formula": {The given target formula if it is not None. The proposed formula in the thought if the thought is correct. Otherwise, choose an appropriate target formula from all possible correct formulas obtained from the tool function for the player to reach. }, "correction": {Json object, the correct thought. None if the thought is correct} } [Current Formula] ... [Thought] ... [Target Formula] ...
@@
Prompt adopted by the VLM corrector model for the ALFWorld task System Prompt: You are an expert in the ALFRED Embodied Environment. The environment requires the player to navigate, take certain objects, interact with objects if necessary, and finally put objects in the designated place to complete the task. Query: You will be given the visual observation and thought of a player in this environment. The task is to ... You are also given the previous actions the player has taken: ... All admissible actions of the current situation are: ... Please evaluate if the reasoning is correct by the following aspects: (1) What objects are in your sight and whether you are holding a certain object? Does the thought correctly identify the image? (2) Based on the task description and the action history, what should be the player’s next sub-goal (notice that the tasks require the player to first pick up certain objects, interact with receptacles if the task is cooling, heating, cleaning or looking in light, and finally placing the object)? Does the thought align with the sub-goal? (3) Based on the task description and the action history, does the player choose one of the admissible actions to reach the sub-goal? Does the action take effect? If the target object is not in sight, go to an unexplored location; if there is a required object, take it; if the task requires cooling, heating, cleaning, or looking in light, navigate and interact with the receptacles. Please briefly answer the above questions, then give your final evaluation. If the thought is incorrect, include all available information for thought correction: choose one correct step from the admissible actions for the player to finish the task, and finally provide the correct thought. Your response should be a valid json file in the following format: { "answer1": {Text, answer to the first question}, "answer2": {Text, answer to the second question}, "answer3": {Text, answer to the third question}, "evaluation": {YES or NO}, "correction": {Json object, the correct thought. None if the thought is correct} } [The player’s thought] ...
Appendix F Example Trajectories of Thought Collapse and GTR
@@
In this section, we present real episode examples from the Points24 and ALFWorld tasks that exhibit thought collapse during RL training. These examples illustrate how the model’s outputs deteriorate, becoming state-irrelevant and irrational, thereby significantly impairing the agent’s decision-making capabilities. This catastrophic phenomenon is the core issue addressed in our work. We also show the GTR-generated trajectories that effectively solve the problem.
F.1 Points24
Figure 15: Examples of agent trajectory produced by RL4VLM (left) and GTR (right) in Points24.
F.2 ALFWorld
Figure 16: Examples of agent trajectory produced by RL4VLM (left) and GTR (right) in ALFWorld.